OpenStudy (anonymous):

Factor -2x^2+12x+14=0

OpenStudy (anonymous):

x=-1,7

sam (.sam.):

-2x^(2)+12x+14=0 Rewrite it 2x^(2)-12x-14=0 Factor 2 out 2(x^(2)-6x-7)=0 Factor the quadratic 2(x+1)(x-7)=0 Ignore the 2, (x+1)(x-7)=0 x=-1,7

OpenStudy (anonymous):

This is a special quadratic. Notice that a-b+c=0 For such quadratic x=-1 and x=-c/a=(-14)/(-2)=7